# Multi-Modal Federated Learning for Cancer Staging over Non-IID Datasets with Unbalanced Modalities

### v1

@@ IV-C Proximity-Aware Client Weighting (PCW)
We propose PCW to tackle the problem of non-iid distribution of local datasets, which leads to the bias of locally trained models in multi-modal FL. PCW defines a measure to weigh the institutions’ local model losses (i.e., $\mathcal{L}_{n}^{\mathsf{Va},(t)}$ and $\mathcal{L}_{n}^{\mathsf{Tr},(t)}$ given in (IV-B) and (IV-B)) to obtain the normalized losses (i.e., ¯$\overline{\mathcal{L}}_{C}^{\mathsf{Va},(t)}$ and ¯$\overline{\mathcal{L}}_{C}^{\mathsf{Tr},(t)}$ used in (11) and (12), and subsequently (13). This measure is defined based on the quality of local datasets of institutions. In order to obtain this measure, we rely on a key observation: considering that the ML models at institutions are initialized with a unique global model at the beginning of each local training round, the difference in their local gradient trajectories is an indicator of the difference in their local data distributions. Furthermore, the difference/closeness between the local gradient trajectories and the gradient trajectory of the global model is an indicator of the similarity of local datasets to the global dataset (i.e., institutions with higher similarity of local gradient trajectory to that of the global model have a closer local data distribution to that of global dataset).
@@
$\Delta\bm{\omega}_{C,G}^{(t)}=[\{\widehat{\bm{\omega}}^{(t)}\}_{m\in C}:\widehat{\bm{\varpi}}_{C}^{(t)}]-[\{\widehat{\bm{\omega}}^{(t+1)}\}_{m\in C}:\widehat{\bm{\varpi}}_{C}^{(t+1)}].$ (17)
@@
Considering (18), $\rho_{C,n}^{(t)}$ takes larger values when local gradient trajectory $\Delta\bm{\omega}_{C,n}^{(t)}$ exhibits a higher similarity to that of global gradient trajectory $\Delta\bm{\omega}_{C}^{(t)}$. [^5] In addition to inner product, we also experimented with other metrics of similarity between the vectors, such as the infimum of the difference in the parameters. Nevertheless, inner product exhibited the best result in our experiments, and thus the rest of the metrics are omitted for brevity. To have comparable values of $\rho_{C,n}^{(t)}$ across institutions $n\in\mathcal{N}$, we normalize the values obtained in (18) to obtain a normalized similarity metric as follows:
$\overline{\rho}_{C,n}^{(t)}={\rho_{C,n}^{(t)}}\Big{/}\Big{(}{\sum_{n\in\mathcal{N}}b_{C,n}\rho_{C,n}^{(t)}}\Big{)},$ (19)
where $b_{C,n}$ is a binary indicator that takes the value of $1$ when where the modalities at institution $n$ match the combination $C$ (i.e., $\mathcal{M}_{n}=C$); otherwise $b_{C,n}=0$.
Having the weights for each institution possessing subset $C$ of modalities, we then derive the average train and validation loss of the respective subset of modalities in the system at the end of each global aggregation round $t$, denoted by ¯$\overline{\mathcal{L}}^{\mathsf{Tr},(t)}_{C}$ and ¯$\overline{\mathcal{L}}^{\mathsf{Va},(t)}_{C}$, respectively, as
@@
Using the estimates in (20), we can substitute the values for ¯$\overline{\mathcal{L}}^{\mathsf{Tr},(t)}_{C}$ and ¯$\overline{\mathcal{L}}^{\mathsf{Va},(t)}_{C}$ in (11) and (12) to calculate the overfitting and generalization, and consequently we can calculate the weights in (13). Compared to a naïve averaging of the training and validation losses, this new formulation of ¯$\overline{\mathcal{L}}^{\mathsf{Tr},(t)}_{C}$ and ¯$\overline{\mathcal{L}}^{\mathsf{Va},(t)}_{C}$ enhances the performance by weighing losses from institutions with closer data distribution to $\mathcal{D}$ with higher values. This is an important step since the average losses being calculated in (11) and (12) will be used for the entire system in (13). Therefore, the institutions where data induces more bias should be given a smaller weight/importance when calculating the average loss, which will determine the DGB weights for all institutions. Our overall methodology for multi-modal FL comprising DGB and PCW is provided in Algorithm 1.
@@ V Numerical Experiments
input : Clients $\mathcal{N}$, Datasets of clients $\{\mathcal{D}_{n}\}_{n\in\mathcal{N}}$, Number of global steps $T$, Number of local SGD iterations $K$ in each global step, Initial weights for gradient blending $\Gamma^{(0)}$
1 for global aggregation round $t=1$ to $T$ do
2       % On the Clients’ Side
3       for client $n=1$ to $|\mathcal{N}|$ in parallel do
@@
5                   Set the encoders’ learning rate coefficients based on $\Gamma^{(0)}$
@@
7                  Calculate the encoder learning rate coefficients $\Gamma_{m},\forall m\in\mathcal{M}_{n}$ based on (13)
@@
9             Send the train loss $\mathcal{L}_{n}^{\mathsf{Tr}}(\bm{\omega}_{n}^{(t),K})$, validation loss $\mathcal{L}_{n}^{\mathsf{Va}}(\bm{\omega}_{n}^{(t),K})$, and the model weights $\bm{\omega}_{n}^{(t),K}$ to the server
@@
12             Aggregate the encoders with weight $\bm{\omega}_{m,n}^{(t),K}$ based on (7).
13      for classifier $c=1$ to $|\mathcal{P}(\mathcal{M})|$ do
14             Aggregate the classifiers with weight $\bm{\varpi}_{n}^{(t),K}$ based on (8).
@@
16             calculate the weight $\overline{\rho}_{m,n}^{(t)}$
@@
18                   calculate the average training and validation losses in global round $t$, ¯$\overline{\mathcal{L}}^{\mathsf{Va},(t)}_{C}$ and ¯$\overline{\mathcal{L}}^{\mathsf{Tr},(t)}_{C}$ based on (20)
19
20
21      Send the calculated $\bm{\omega}_{m,n}^{(t+1),0}$, $\bm{\varpi}_{C}^{(t+1),0}$, {¯$\{\overline{\mathcal{L}}^{\mathsf{Tr},(t)}_{C}\}_{C\in\mathcal{P}(\mathcal{M})}$, and {¯$\{\overline{\mathcal{L}}^{\mathsf{Va},(t)}_{C}\}_{C\in\mathcal{P}(\mathcal{M})}$ back to the clients.
@@ V-A Dataset and Data Distribution
We use datasets from The Cancer Genome Atlas program (TCGA) for three cancer types, namely, Breast Invasive Carcinoma (BRCA), Lung Squamous Cell Carcinoma (LUSC), and Liver Hepatocellular Carcinoma (LIHC), to make up for the heterogeneity of local datasets across clients. Each dataset has three modalities: mRNA sequence, histopathological images, and clinical information. Only the patient data that have all three modalities to make the results comparable to centralized ML as conducted in [9] and the FL settings described in [31, 32]. Nevertheless, in our setting, as will be described later, the data is distributed across the system such that each institution has access to only a fraction of the data modalities. The cancer stage distribution for the three datasets is described in Table I.
TABLE I: Data distribution across cancer stages in the datasets used.
Stage Type
BRCA
LUSC
LIHC
Stage 1
190
66
51
Stage 2
55
45
39
@@ V-C Baseline Methods
    UM-FL: The upper bound FL performance is achieved when all institutions possess all of data modalities, removing the barrier of different convergence rates across modalities. This baseline is only added to provide an estimate of the achievable performance.
@@
We consider the first baseline to measure the system’s performance under the conventional methods (i.e., FedAvg aggregation and local training using SGD), which is the naïve extension of FL to multi-modal setting as described in [32]. We will demonstrate that this approach will perform poorly over unbalanced modalities across the institutions. The second baseline [31] implements a federated version of gradient blending introduced in [21], which is not tailored to address the problem of uneven number of modalities across institutions. This baseline also relies on tuning the global model aggregations weights, which is different from our method. Further, our method removes the necessity to split the training data at each institution, as done in [21], to calculate the gradient blending weights. This is reflected in our formulation (20) where we aggregate all of the loss values needed to calculate the DGB parameters based on the modality combination of the institutions to get a distributed calculation of these parameters. Avoiding splitting the dataset in data-scarce scenarios such as ours, and generally, most medical applications, can (later shown in Table III) improve the outcome of the federated learning. We further complement our method with PCW, which as will be shown later will give us performance edge. In the third baseline, we aim to measure the base performance of our proposed DGB and demonstrate the performance gain we get by employing the method. Although we use this method as a baseline, we will later show that this method outperforms the state-of-the-art, e.g., [32, 31], due to consideration of different convergence speeds across various modalities.

### v2

@@ IV-C Proximity-Aware Client Weighting (PCW)
We propose PCW to tackle the problem of non-iid distribution of local datasets, which leads to the bias of locally trained models in multi-modal FL. PCW defines a measure to weigh the institutions’ local model losses (i.e., $\mathcal{L}_{n}^{\mathsf{Va},(t)}$ and $\mathcal{L}_{n}^{\mathsf{Tr},(t)}$ given in (IV-B) and (IV-B)) to obtain the normalized losses (i.e., ¯$\overline{\mathcal{L}}_{C}^{\mathsf{Va},(t)}$ and ¯$\overline{\mathcal{L}}_{C}^{\mathsf{Tr},(t)}$ used in (11) and (12), and subsequently (13). This measure is defined based on the quality of local datasets of institutions. In order to obtain this measure, we rely on a key observation: considering that the ML models at institutions are initialized with a unique global model at the beginning of each local training round, the difference in their local gradient trajectories (i.e., their cumulative gradient generated during their local ML model training) is an indicator of the difference in their local data distributions. Furthermore, the difference/closeness between the local gradient trajectories and the gradient trajectory of the global model is an indicator of the similarity of local datasets to the global dataset (i.e., institutions with higher similarity of local gradient trajectory to that of the global model have a closer local data distribution to that of global dataset, and thus a higher data quality). Using this measure, we will be able to calculate ¯$\overline{\mathcal{L}}_{C}^{\mathsf{Va},(t)}$ and ¯$\overline{\mathcal{L}}_{C}^{\mathsf{Tr},(t)}$, while explicitly considering the non-iid nature of data across the institutions. In particular, ¯$\overline{\mathcal{L}}_{C}^{\mathsf{Va},(t)}$ and ¯$\overline{\mathcal{L}}_{C}^{\mathsf{Tr},(t)}$ are later computed through assigning a smaller weight to the loss of institutions with more biased (i.e., lower quality) datasets, while assigning a higher weight to the loss of institutions with higher quality data. More reliance/emphasis on the loss values of institutions with higher data quality through PCW, leads to a better capture of the behavior of each specific modality combination for the global distribution of data.
@@
$\Delta\bm{\omega}_{C,G}^{(t)}=\Big{[}\{\widehat{\bm{\omega}}_{m}^{(t)}\}_{m\in C}:\widehat{\bm{\varpi}}_{C}^{(t)}\Big{]}-\Big{[}\{\widehat{\bm{\omega}}^{(t+1)}_{m}\}_{m\in C}:\widehat{\bm{\varpi}}_{C}^{(t+1)}\Big{]}.$ (17)
@@
Considering (18), $\rho_{C,n}^{(t)}$ takes larger values when local gradient trajectory $\Delta\bm{\omega}_{C,n}^{(t)}$ exhibits a higher similarity to that of global gradient trajectory $\Delta\bm{\omega}_{C}^{(t)}$. [^10] In addition to inner product, we also experimented with other metrics of similarity between the vectors, such as the infimum of the difference in the parameters. Nevertheless, inner product exhibited the best result in our experiments, and thus the rest of the metrics are omitted for brevity. To have comparable values of $\rho_{C,n}^{(t)}$ across institutions, we normalize the values obtained in (18) to obtain a normalized positive similarity metric as follows:
$\overline{\rho}_{C,n}^{(t)}={e^{\tau\rho_{C,n}^{(t)}}}\Big{/}\Big{(}{\sum_{n^{\prime}\in\mathcal{N}}b_{C,n^{\prime}}e^{\tau\rho_{C,n^{\prime}}^{(t)}}}\Big{)},$ (19)
where $b_{C,n^{\prime}}$ follows the same definition as in (8) and $\tau$ is a tunable temperature (chosen as $\tau=1$ in our simulations).
Having the train and validation losses, i.e., (IV-B) and (IV-B), for each institution possessing subset $C$ of modalities, we then derive the adjusted average train and validation loss of the respective subset of modalities in the system at the end of each global aggregation step $t$, denoted by ¯$\overline{\mathcal{L}}^{\mathsf{Tr},(t)}_{C}$ and ¯$\overline{\mathcal{L}}^{\mathsf{Va},(t)}_{C}$, respectively, via the integration of the above similarity metric as follows ($b_{C,n}$ is defined the same as in (8)):
@@
Using the estimates in (20), we can substitute the values for ¯$\overline{\mathcal{L}}^{\mathsf{Tr},(t)}_{C}$ and ¯$\overline{\mathcal{L}}^{\mathsf{Va},(t)}_{C}$ in (11) and (12) to calculate the overfitting and generalization, and consequently we can calculate the weights in (13). Compared to a naïve averaging of the training and validation losses, this new formulation of ¯$\overline{\mathcal{L}}^{\mathsf{Tr},(t)}_{C}$ and ¯$\overline{\mathcal{L}}^{\mathsf{Va},(t)}_{C}$ enhances the performance by weighing losses from institutions with closer data distribution to the global data distribution $\mathcal{D}$ with higher values, thus tackling the challenge of non-iid distributed data when calculating the DOGR coefficients. This is an important step since the average losses being calculated in (11) and (12) will be used for the entire system in (13). Therefore, the institutions where data induces more bias should be given a smaller weight/importance when calculating the average loss, which will determine the DGB weights for all institutions. Our overall methodology for multi-modal FL comprising DGB and PCW is provided in Algorithm 1.
@@ V Numerical Experiments
input : Institutions $\mathcal{N}$, Datasets of institutions $\{\mathcal{D}_{n}\}_{n\in\mathcal{N}}$, Number of global steps $T$, Number of local SGD iterations $K$ in each global step, Initial DOGR weights for modality encoders $\Gamma^{(0)}_{m,n}$, Initial DOGR weights for institution classifiers $\Gamma^{(0)}_{\mathcal{M}_{n},n}$
1 for global aggregation step $t=1$ to $T$ do
2       % On the institutions’ Side
3       for institution $n=1$ to $|\mathcal{N}|$ in parallel do
@@
5                   Set encoder of modality $m$’s learning rate coefficients $\Gamma^{(t)}_{m,n}=\Gamma^{(0)}_{m,n}$
6                   Set the classifier’s learning rate coefficients $\Gamma^{(t)}_{\mathcal{M}_{n},n}=\Gamma^{(0)}_{\mathcal{M}_{n},n}$
@@ 5                   Set encoder of modality $m$’s learning rate coefficients $\Gamma^{(t)}_{m,n}=\Gamma^{(0)}_{m,n}$
8                  Calculate the encoder and classifier learning rate coefficients $\Gamma_{m,n}^{(t)},\forall m\in\mathcal{M}_{n}$ and $\Gamma_{\mathcal{M}_{n},n}^{(t)}$ using (13)
@@
10             Send the train loss $\mathcal{L}_{n}^{\mathsf{Tr}}(\bm{\omega}_{n}^{(t),K})$, validation loss $\mathcal{L}_{n}^{\mathsf{Va}}(\bm{\omega}_{n}^{(t),K})$, and the model parameters $\bm{\omega}_{n}^{(t),K}=\big{\{}\{\bm{\omega}_{m,n}^{(t),K}\}_{m\in\mathcal{M}_{n}}:\bm{\varpi}^{(t),K}_{n}\big{\}}$ to the server
@@
13             Aggregate the encoders with parameters $\bm{\omega}_{m,n}^{(t),K}$ based on (7) to obtain $\widehat{\bm{\omega}}_{m}^{(t+1)}$.
14      for combination $C\in\mathcal{P}(\mathcal{M})$ do
15             Aggregate the classifiers with parameters $\bm{\varpi}_{n}^{(t),K}$ based on (8) to obtain $\widehat{\bm{\varpi}}_{C}^{(t+1)}$.
@@
17             Calculate the weight $\overline{\rho}_{C,n}^{(t)}$ based on (19)
@@
19                   Calculate the average training and validation losses in global aggregation step $t$, ¯$\overline{\mathcal{L}}^{\mathsf{Va},(t)}_{C}$ and ¯$\overline{\mathcal{L}}^{\mathsf{Tr},(t)}_{C}$ based on (20)
20                   Calculate $O_{C}^{(t)}$ and $G_{C}^{(t)}$ based on (11) and (12)
21
22      Send the calculated $\widehat{\bm{\omega}}_{m}^{(t+1)}$, $\widehat{\bm{\varpi}}_{C}^{(t+1)}$, $O_{C}^{(t)}$ and $G_{C}^{(t)},C\in\mathcal{P}(\mathcal{M})$ back to the institutions.
@@ V-A Dataset and Data Distribution
We use datasets from The Cancer Genome Atlas program (TCGA) [35] for three cancer types, namely, Breast Invasive Carcinoma (BRCA), Lung Squamous Cell Carcinoma (LUSC), and Liver Hepatocellular Carcinoma (LIHC). From each cohort, we will use three of the available modalities: mRNA sequence, histopathological images, and clinical information. The data from these cohorts is filtered to keep only those patient data points that have all three modalities to make the results comparable to centralized ML as conducted in [9] and the FL settings described in [31, 32]. We use this data to simulate each institution’s local dataset in our experiments. The distribution of the class labels across the cohorts is summarized in Table II.
TABLE II: Available patients for each stage across the cohorts used for binary and multi-class classification.
Stage Cohort BRCA LUSC LIHC All Cohorts (Binary/Multi-Class) Stage Ratio (%) (Binary/Multi-Class)
Stage I 155 152 152 459 $\simeq$40.55/31.48
Stage II 488 108 77 673 $\simeq$59.45/46.16
[…]
All Stages (Binary/Multi-Class) 643/851 260/304 229/303 1132/1458
Cohort Ratio (%) (Binary/Multi-Class) $\simeq$56.80/58.37 $\simeq$22.97/20.85 $\simeq$20.23/20.78
We next distribute the refined dataset across $21$ institutions, where, to emulate real-world settings, we consider the existence of unbalanced modality combinations across the institutions. In particular, the configuration of modalities in each institution’s local dataset is summarized in Table III. Given both binary classification [2, 5, 7, 9] and multi-class classification [3, 4] are considered in the existing literature on the task of cancer staging, to provide benchmarks for the follow-up works in the emerging area of multi-modal FL, we consider both binary and multi-class classification and present their respective results in this section.
Next, we further delve into the reflection of non-iid-ness/heterogeneity of data across the institutions. In particular, in our network of interest with multiple data modalities and various stages of cancer (i.e., data labels), non-iid-ness/heterogeneity of data can be interpreted at two levels: (i) type-based heterogeneity, and (ii) class-based heterogeneity, as elaborated below.
TABLE III: Modality combinations used across institutions.
Data Modalities Entity
mRNA Image Clinical Institution #
1, 2, 3
4, 5, 6
7, 8, 9
10, 11, 12
13, 14, 15
16, 17, 18
19, 20, 21
  • Type-based Heterogeneity: We keep the same ratio of class labels across institutions’ local datasets; $\mathcal{D}_{n}\leavevmode\nobreak\ \forall n$ contain the same fraction of data points of different labels. However, the ratio of the data points belonging to various cohorts will be different across institutions, which leads to a bias toward a specific cancer type in each institution’s local data. Following the modality spread across the institutions given in Table III, the fraction of each cohort’s data inherited by each institution is given in Table IV for each of the binary and multi-class classification scenarios.
  • Class-based Heterogeneity: Institutions’ local datasets will have the same ratio of datapoints of the three cancer types. However, institutions’ local datasets contain a varying ratio of datapoints belonging to different classes (i.e., cancer stages), which leads to a bias toward a specific cancer stage/class in each institution’s local data. The class distribution of data for binary and multi-class classification cases is shown in Table V.
TABLE IV: Type-based heterogeneity data distribution across institutions based on the fraction of institutions’ local dataset consisting of each cohort’s data.
Class Labels (Binary Classification) Class Labels (Multi-Class Classification) Entity
BRCA LUSC LIHC BRCA LUSC LIHC Institution # Category
$\simeq$ 56% $\simeq$ 22% $\simeq$ 20% $\simeq$ 58% $\simeq$ 21% $\simeq$ 21% 1, 4, 7, 10, 13, 16, 19 1
$\simeq$ 54% $\simeq$ 10% $\simeq$ 35% $\simeq$ 48% $\simeq$ 28% $\simeq$ 24% 2, 5, 8, 11, 14, 17, 20 2
$\simeq$ 58% $\simeq$ 35% $\simeq$ 6% $\simeq$ 62% $\simeq$ 12% $\simeq$ 26% 3, 6, 9, 12, 15, 18, 21 3
TABLE V: Class-based heterogeneity data distribution across institutions based on the fraction of institutions’ local dataset consisting of each class label.
Class Labels (Binary Classification) Class Labels (Multi-Class Classification) Entity
Stage I Stage II Stage I Stage II Stage III Institution # Category
$\simeq$ 41% $\simeq$ 59% $\simeq$ 32% $\simeq$ 46% $\simeq$ 22% 1, 4, 7, 10, 13, 16, 19 1
$\simeq$ 62% $\simeq$ 38% $\simeq$ 36% $\simeq$ 34% $\simeq$ 30% 2, 5, 8, 11, 14, 17, 20 2
$\simeq$ 19% $\simeq$ 81% $\simeq$ 30% $\simeq$ 54% $\simeq$ 16% 3, 6, 9, 12, 15, 18, 21 3
Heterogeneity Severity Comparison. It is worth mentioning that type-based heterogeneity can be though of as a more fundamental level of heterogeneity in which the institutions local datasets have unbalanced ratios of the cancer types as compared to the unbalanced ratio of their labels imposed in class-based heterogeneity. As a result, type-based heterogeneity will have a more pronounced impact on the model training since during the local model training at each institution, for each cancer type, a specific part of the input information [^11] For example, specific elements in the mRNA sequence, specific extracted image features, specific clinical information is important: having unbalanced ratios of cancer types may lead to the confusion of the local model regarding the identification of the most important parts of the input information to focus for conducting classification. Nevertheless, since the decisive input features for classification may be shared across different classes for the same cancer type, imposing class-based heterogeneity while having the same distribution of cancer types across institutions will not result in such a severe bias in the local ML models during training. For example, to differentiate between different stages of a cancer the model may focus on the same feature set of the histopathological images, which may correspond to the location of a sensitive tissue. Nevertheless, such feature set (or the location of focus) may differ across different cancers. Thus, having unbalanced ratios of cancer types may bias the model towards focusing on the input features of the majority type of the local datapoints, which may differ across cancer types; however, having unbalanced classes may not impact the local model bias as much.
@@ V-C Baseline Methods
    UM-FL: The upper bound FL performance is achieved when all institutions possess all of data modalities (i.e., the scenario resembles uniform-modality distribution), removing the barrier of different convergence rates across modalities. This baseline is only added to provide an estimate of the achievable performance.
TABLE VI: Different architectures for the mRNA encoder.
Layer Type Model Depth
Small Medium Large
Input Layer FC (20531, ReLU)
Layer 1 FC (4096, ReLU) FC (8192, ReLU) FC (16384, ReLU)
Layer 2 FC (2048, ReLU) FC (4096, ReLU) FC (8192, ReLU)
Layer 3 FC (512, ReLU) FC (2048, ReLU) FC (4096, ReLU)
Layer 4 FC (128, ReLU) FC (512, ReLU) FC (2048, ReLU)
Layer 5 FC (64, ReLU) FC (128, ReLU) FC (512, ReLU)
Layer 6 FC (32, ReLU) FC (64, ReLU) FC (128, ReLU)
Layer 7 FC (32, ReLU) FC (64, ReLU)
Layer 8 FC (32, ReLU)
Output Layer Binary/Multi-Class FC (2, LogSoftmax) / FC (3, LogSoftmax)
TABLE VII: Accuracy (%) of different methods throughout the global aggregations under both binary and multi-class classification for type-based heterogeneity.
Binary Classification Multi-Class Classification
Global Round Method CM-FL HGB-FL CM-FL + DGB Our Method UM-FL (Upperbound) Global Round Method CM-FL HGB-FL CM-FL + DGB Our Method UM-FL (Upperbound)
0 $\simeq$ 52.34 $\simeq$ 52.34 $\simeq$ 52.34 $\simeq$ 52.34 $\simeq$ 52.34 0 $\simeq$ 36.52 $\simeq$ 36.52 $\simeq$ 36.52 $\simeq$ 36.52 $\simeq$ 36.52
100 $\simeq$ 62.92 $\simeq$ 62.39 $\simeq$ 63.34 $\simeq$ 61.80 $\simeq$ 69.32 100 $\simeq$ 41.69 $\simeq$ 49.32 $\simeq$ 46.26 $\simeq$ 49.93 $\simeq$ 54.74
200 $\simeq$ 59.55 $\simeq$ 63.25 $\simeq$ 63.64 $\simeq$ 64.04 $\simeq$ 73.38 200 $\simeq$ 44.26 $\simeq$ 55.14 $\simeq$ 48.25 $\simeq$ 58.63 $\simeq$ 63.18
300 $\simeq$ 62.92 $\simeq$ 64.10 $\simeq$ 64.24 $\simeq$ 68.54 $\simeq$ 76.24 300 $\simeq$ 46.58 $\simeq$ 56.17 $\simeq$ 49.63 $\simeq$ 64.32 $\simeq$ 68.25
400 $\simeq$ 61.80 $\simeq$ 65.81 $\simeq$ 66.43 $\simeq$ 71.91 $\simeq$ 79.46 400 $\simeq$ 47.52 $\simeq$ 56.83 $\simeq$ 52.36 $\simeq$ 66.24 $\simeq$ 72.85
500 $\simeq$ 64.04 $\simeq$ 67.52 $\simeq$ 67.13 $\simeq$ 74.16 $\simeq$ 79.46 500 $\simeq$ 47.82 $\simeq$ 57.14 $\simeq$ 53.64 $\simeq$ 67.35 $\simeq$ 74.47
600 $\simeq$ 61.80 $\simeq$ 68.64 $\simeq$ 69.21 $\simeq$ 74.16 $\simeq$ 80.24 600 $\simeq$ 49.49 $\simeq$ 57.52 $\simeq$ 53.83 $\simeq$ 67.84 $\simeq$ 75.36
TABLE VIII: Accuracy (%) of different methods throughout the global aggregations under both binary and multi-class classification for class-based heterogeneity.
Binary Classification Multi-Class Classification
Global Round Method CM-FL HGB-FL CM-FL + DGB Our Method UM-FL (Upperbound) Global Round Method CM-FL HGB-FL CM-FL + DGB Our Method UM-FL (Upperbound)
0 $\simeq$ 52.34 $\simeq$ 52.34 $\simeq$ 52.34 $\simeq$ 52.34 $\simeq$ 52.34 0 $\simeq$ 36.52 $\simeq$ 36.52 $\simeq$ 36.52 $\simeq$ 36.52 $\simeq$ 36.52
100 $\simeq$ 53.58 $\simeq$ 63.14 $\simeq$ 65.46 $\simeq$ 63.48 $\simeq$ 70.58 100 $\simeq$ 43.14 $\simeq$ 51.37 $\simeq$ 46.83 $\simeq$ 55.32 $\simeq$ 56.82
200 $\simeq$ 57.13 $\simeq$ 64.35 $\simeq$ 67.18 $\simeq$ 66.80 $\simeq$ 75.23 200 $\simeq$ 47.23 $\simeq$ 56.29 $\simeq$ 53.82 $\simeq$ 60.82 $\simeq$ 63.43
300 $\simeq$ 62.43 $\simeq$ 64.66 $\simeq$ 68.12 $\simeq$ 68.23 $\simeq$ 78.45 300 $\simeq$ 47.23 $\simeq$ 58.34 $\simeq$ 55.72 $\simeq$ 64.17 $\simeq$ 69.72
400 $\simeq$ 65.24 $\simeq$ 66.73 $\simeq$ 69.57 $\simeq$ 73.28 $\simeq$ 79.75 400 $\simeq$ 47.82 $\simeq$ 62.04 $\simeq$ 51.93 $\simeq$ 66.82 $\simeq$ 72.55
500 $\simeq$ 64.04 $\simeq$ 68.65 $\simeq$ 70.43 $\simeq$ 76.24 $\simeq$ 80.34 500 $\simeq$ 48.53 $\simeq$ 63.18 $\simeq$ 56.72 $\simeq$ 68.39 $\simeq$ 74.28
600 $\simeq$ 62.84 $\simeq$ 69.15 $\simeq$ 71.13 $\simeq$ 76.31 $\simeq$ 81.26 600 $\simeq$ 48.53 $\simeq$ 63.56 $\simeq$ 57.26 $\simeq$ 69.92 $\simeq$ 75.10
TABLE IX: Global model accuracy (%) for each dominant modality scenario for both binary classification and multi-class classification cases in the presence of type-based heterogeneity in local institution datasets.
Binary Classification Multi-Class Classification
Method Dominant Modality Uni-Modal Bi-Modal Tri-Modal Method Dominant Modality Uni-Modal Bi-Modal Tri-Modal
CM-FL $\simeq$ 60.12 $\simeq$ 60.56 $\simeq$ 62.25 CM-FL $\simeq$ 46.59 $\simeq$ 47.62 $\simeq$ 48.85
CM-FL + DGB $\simeq$ 67.14 $\simeq$ 68.23 $\simeq$ 70.83 CM-FL + DGB $\simeq$ 54.27 $\simeq$ 55.61 $\simeq$ 58.25
Our Method $\simeq$ 72.75 $\simeq$ 73.21 $\simeq$ 74.82 Our Method $\simeq$ 66.43 $\simeq$ 67.82 $\simeq$ 69.17
TABLE X: Global model accuracy (%) for each dominant modality scenario for both binary classification and multi-class classification cases in the presence of class-based heterogeneity in local institution datasets.
Binary Classification Multi-Class Classification
Method Dominant Modality Uni-Modal Bi-Modal Tri-Modal Method Dominant Modality Uni-Modal Bi-Modal Tri-Modal
CM-FL $\simeq$ 62.34 $\simeq$ 63.02 $\simeq$ 65.42 CM-FL $\simeq$ 49.75 $\simeq$ 51.09 $\simeq$ 53.29
CM-FL + DGB $\simeq$ 67.43 $\simeq$ 68.65 $\simeq$ 71.20 CM-FL + DGB $\simeq$ 60.13 $\simeq$ 64.14 $\simeq$ 67.84
Our Method $\simeq$ 75.30 $\simeq$ 75.83 $\simeq$ 76.63 Our Method $\simeq$ 67.17 $\simeq$ 68.18 $\simeq$ 70.32
@@
We consider the first baseline to measure the system’s performance under the conventional methods (i.e., FedAvg aggregation and local training using SGD), which is the naïve extension of FL to multi-modal setting as described in [32]. We will demonstrate that this approach will perform poorly over unbalanced modalities across the institutions. The second baseline [31] implements a federated version of gradient blending introduced in [21], which is not tailored to address the problem of uneven number of modalities across institutions. This baseline also relies on tuning the global model aggregations weights, which is different from our method. Further, our method removes the necessity to split the training data at each institution, as done in [21], to calculate the gradient blending weights. This is reflected in our formulation (20) where we aggregate all of the loss values needed to calculate the DGB parameters based on the modality combination of the institutions to get a distributed calculation of these parameters. Avoiding splitting the dataset in data-scarce scenarios such as ours, and generally, most medical applications, can (later shown in Sec. V-E1) improve the outcome of the federated learning. We further complement our method with PCW, which as will be shown later will give us performance edge. In the third baseline, we aim to measure the base performance of our proposed DGB and demonstrate the performance gain we get by employing the method. Although we use this method as a baseline, we will later show that this method outperforms the state-of-the-art, e.g., [32, 31], due to consideration of different convergence speeds across various modalities.